### v1

# Measuring What Makes You Unique: Difference-Aware User Modeling for Enhancing LLM Personalization

Abstract
Personalizing Large Language Models (LLMs) has become a critical step in facilitating their widespread application to enhance individual life experiences. In pursuit of personalization, distilling key preference information from an individual’s historical data as instructional preference context to customize LLM generation has emerged as a promising direction. However, these methods face a fundamental limitation by overlooking the inter-user comparative analysis, which is essential for identifying the inter-user differences that truly shape preferences. To address this limitation, we propose Difference-aware Personalization Learning (DPL), a novel approach that emphasizes extracting inter-user differences to enhance LLM personalization. DPL strategically selects representative users for comparison and establishes a structured standard to extract meaningful, task-relevant differences for customizing LLM generation. Extensive experiments on real-world datasets demonstrate that DPL significantly enhances LLM personalization. We release our code at https://github.com/SnowCharmQ/DPL.
Measuring What Makes You Unique: Difference-Aware User Modeling for Enhancing LLM Personalization
Yilun Qiu1††thanks: Equal contribution, Xiaoyan Zhao211footnotemark: 1, Yang Zhang1††thanks: Corresponding Author, Yimeng Bai3, Wenjie Wang3, Hong Cheng2, Fuli Feng3, Tat-Seng Chua1, 1National University of Singapore 2The Chinese University of Hong Kong 3University of Science and Technology of China qiuyilun@u.nus.edu, xzhao@se.cuhk.edu.hk, zyang1580@gmail.com, baiyimeng@mail.ustc.edu.cn, wenjiewang96@gmail.com, hcheng@se.cuhk.edu.hk, fulifeng93@gmail.com, dcscts@nus.edu.sg
1 Introduction
With continuous efforts, Large Language Models (LLMs) Achiam et al. (2023); Yang et al. (2024); Guo et al. (2025); Dubey et al. (2024) have demonstrated increasingly higher levels of intelligence, sparking an unprecedented enthusiasm for applying them to individuals’ daily lives to enhance personal thinking, planning, and overall life experiences Bao et al. (2023); Razafinirina et al. (2024); Li et al. (2024b); Deng et al. (2024); Zhang et al. (2025b). However, the “one-size-fits-all” paradigm of general LLMs becomes ineffective when serving individual users, as users have distinct preferences and would expect personalized generation from LLMs Christakopoulou et al. (2023). This motivates an exciting new direction: LLM personalization Zhang et al. (2024b) – adapting LLMs to generate text responses that align with each individual’s unique preferences.
LLM personalization fundamentally relies on incorporating user-specific information into LLMs to tailor their generation Kang et al. (2023); Li et al. (2023b). The “memorize-then-inject" framework offers a promising and economical solution – storing users’ historical data and then extracting key information to serve as instructional contexts for customizing LLM responses Salemi et al. (2024b). Depending on the extraction method employed, existing approaches can generally be categorized into: 1) Retrieval-based methods Mysore et al. (2024); Li et al. (2023a), which just retrieve the parts of the user’s history relevant to the current request for LLM use, and 2) Summarization-based methods Richardson et al. (2023); Zheng et al. (2024), which further condense the user history to distill preferences for LLM use.
While these methods Mysore et al. (2024); Zheng et al. (2024) have made notable progress, we contend that they share a common limitation: they overlook the inter-user comparative analysis, which is essential for identifying the differences that truly shape preferences. As acknowledged in psychology and behavioral science Snyder and Fromkin (1977, 2012); Irmak et al. (2010), it is the differences among individuals that make them unique, while their uniqueness shapes their preferences. This underscores that effective personalization hinges on identifying and understanding these differences. As a result, without considering inter-user comparisons to extract differences, current methods would be suboptimal for personalization.
In this work, we propose difference-aware user modeling – extracting the differences among users as key preference information – to enhance LLM personalization. To achieve this, distinguishing each user from others and then summarizing the differences (using an LLM) to serve as the instructional contexts for personalization provides an intuitive solution. However, it faces two challenges: 1) comparing to all users may be impractical due to the high cost and complexity of extracting differences across all others; and 2) without proper guidance, the differences extracted by the LLM may be irrelevant or meaningless for the task. Therefore, we must address: 1) how to select appropriate users for comparison, and 2) how to extract task-relevant differences during comparison effectively.
To address these challenges, we introduce a novel Difference-aware Personalization Learning (DPL) method. DPL selectively chooses representative users for comparison and establishes an extraction standard to effectively capture task-relevant differences, using the extracted differences as the LLM’s instructional contexts for personalization. Specifically, we cluster users based on data similarity and select only cluster-center users for comparison, ensuring efficiency while maintaining a global perspective. To extract task-relevant differences, we structure the difference extraction model (an LLM) to focus on key predefined dimensions for capturing differences, rather than operating freely. Three key preference dimensions are distilled from existing literature as examples: writing style, emotional tone, and semantic content.
We validate our DPL method on a representative personalized text generation task, namely, review generation Peng et al. (2024). Extensive results demonstrate that DPL effectively enhances LLM personalization.
The main contributions of this work are summarized as follows:
We emphasize the philosophy that “differences make us unique” as the key to personalization and propose using difference-aware user modeling to enhance LLM personalization.
We propose a novel Difference-aware Personalization Learning method for LLM, incorporating a selective comparison mechanism and structured difference extraction mechanism to achieve effective difference extraction.
Extensive experiments on real-world datasets demonstrate that DPL achieves state-of-the-art performance, highlighting its effectiveness in personalized text generation.
2 Preliminary
Problem Formulation. This work focuses on personalizing LLMs to generate text outputs that align with user preferences, i.e., achieving personalized text generation. We assume that users have a set of historical texts, either written or preferred by them, reflecting the users’ preferences. We could leverage this data to customize LLM’s generation to meet user needs. Formally, let denote the historical data for all users. Each sample in is represented as , where represents the user, denotes the item (or object) the user is focused on, and indicates the text written or preferred by for item . When a target user submits a new request for text generation on a target item , the LLM needs to generate text that satisfies the user’s preferences based on .
Notably, without loss of generality, this paper primarily focuses on review generation, a representative personalized text generation task. In this context, the goal is to enable LLMs to generate personalized reviews for items such as movies or products, ensuring alignment with true user reviews.
Memorize-then-Inject Solution. A typical approach in existing works for customizing LLM generation is to store the user’s historical data and then extract key information from the data to distill preference information through retrieval or summarization when needed. This extracted information serves as a preference context in the instruction to guide the model’s generation. Formally, to generate personalized text for the focused user on the target item , the method can be represented as follows:
where denotes the generated result, denotes the historical data for , and represents the key information extraction process from . Notably, the extraction process usually also involves the utilization of an LLM. This method does not require LLM retraining and can potentially manage lifelong historical data in a memory manner Zhang et al. (2024a), making it a promising and cost-effective solution.
3 Methodology
In this section, we introduce the proposed Difference-Aware Personalization Learning (DPL) method, beginning with its motivation and the overall framework, followed by a detailed explanation of each component.
3.1 Overview
Given that the differences define individual uniqueness, we argue that inter-user differences play a crucial role in preference learning for LLM personalization. However, existing approaches overlook the inter-user comparisons for preference modeling, limiting the ability to extract meaningful differences for personalization. To address this limitation, we propose the DPL method. As shown in Figure 1, DPL introduces an inter-user comparison mechanism to identify differences between users, which are then combined with the user’s own key history to form the preference context in instructions, guiding LLM generation. This process can be formulated as:
where represents the difference-aware preference extracted by comparing a user’s historical data with other users’ data in .
Difference-Aware Preference Extraction Process. As illustrated in the bottom part of Figure 1, for user , we assume that contains key elements, denoted as , which can be obtained via retrieval Robertson and Zaragoza (2009). Then we would compare each element with those of some other users to extract differences and then summarize these differences across all samples and the key historical data to obtain the final preference. Specifically, for each element , we compare it with the texts of other users on the same item , represented by , using a selective user comparison and structured difference extraction process. This process consists of two key steps, as shown in the upper part of the figure:
Selective user comparison: Instead of comparing with all users in , we selectively compare with a subset of representative users to reduce complexity and improve efficiency.
Structured difference extraction: Differences are extracted according to predefined standards to ensure task relevance and meaningfulness.
The details of these steps will be elaborated later. After extracting differences for all elements, we summarize them as well as the key history with an LLM to obtain the final difference-aware representation for personalization. Formally,
where denotes the summarizer, and represents the extracted difference for the -th element in . Notably, we also take into account the user’s own data to retain preference information that cannot be captured by the differences, i.e., using the key history as additional inputs. Next, we elaborate on the selective user comparison and structured difference extraction methods to explain how each is obtained.
3.2 Selective User Comparison and Structured Difference Extraction
When dealing with each important historical data sample , a straightforward approach for difference extraction is to have the LLM compare all users and summarize the differences. However, comparing all users is complex and inefficient. To address this, we selectively compare only representative users. Additionally, since differences can span multiple dimensions, allowing LLMs to extract them freely may introduce task-irrelevant information. To mitigate this, we introduce a structured difference extraction approach with a predefined standard to guide the extraction of meaningful and task-relevant differences.
3.2.1 Selective User Comparison
To select representative users for comparison, the core consideration is that they could represent the total population, making the comparison still provide a global view. To achieve this, we consider clustering the users based on their text and then selecting the user in the cluster center for comparison. Specifically, when dealing with each key element in history for the target , we just consider the users having texts on , and then cluster these users into groups according to the texts corresponding to them, obtaining the most central users for each group. We then select the central users that do not belong to groups of the target user as the users for comparison. Formally,
where denotes the selected representative users, and denotes all the historical data that the users have texts on item . We represent the data samples regarding the selected users on item as: .
3.2.2 Structured Difference Extraction
After obtaining the representative users, we compare their texts with the target user’s texts, i.e., comparing with to obtain the difference via an LLM. Instead of allowing LLMs to freely summarize the difference, we would define a standard to structure the LLM difference summarization. The process can be formulated as:
where denotes the LLM to extract differences. Here, represents the standard that the LLMs follow to generate the difference. Specifically, we distill three key dimensions from existing literature (as examples): writing style, emotional style, and semantic style, and then structure the LLM to extract differences from these dimensions. The three dimensions are detailed below:
Writing style. This dimension captures the lexical choices, syntactic structures, and linguistic patterns unique to each user. By analyzing vocabulary richness, sentence complexity, and grammatical preferences, we can identify how a user’s writing diverges from others Verma and Srinivasan (2019).
Emotional style. This feature reflects the sentiment and affective tone embedded in user’s text. By assessing the polarity (positive, negative, neutral), we can highlight differences in how users convey their attitudes and feelings Kheiri and Karimi (2023).
Semantic style. This aspect focuses on the depth, clarity, and coherence of meaning in user’s text. By evaluating information density, logical flow, and contextual relevance, we can differentiate how users structure and convey their ideas Sahu and Majumdar (2017).
Extracting differences along these dimensions ensures meaningful and well-structured results for distinguishing task-relevant content. We guide the LLM to adhere to these dimensions for difference extraction using prompts. All prompts used in DPL are provided in Appendix F.
4 Experiments
In this section, we conduct experiments to answer the following research questions:
RQ1: How does DPL perform on the personalized text generation task with real-world datasets, in comparison to other baseline methods?
RQ2: What is the impact of the individual components of DPL on its effectiveness?
RQ3: How do the specific hyper-parameters of DPL influence its performance?
RQ4: How does DPL perform across different levels of user uniqueness?
RQ5: How does DPL qualitatively demonstrate its personalized generation capabilities through concrete case analysis?
| Datasets () | Books | Movies & TV | CDs & Vinyl | |||||||||
| Methods () | R-1 | MET. | S-72B | S-GPT | R-1 | MET. | S-72B | S-GPT | R-1 | MET. | S-72B | S-GPT |
| Non-Perso | 0.2683 | 0.1512 | 0.5079 | 0.4360 | 0.2275 | 0.1304 | 0.4767 | 0.3242 | 0.2392 | 0.1336 | 0.4701 | 0.3409 |
| RAG | 0.3206 | 0.2173 | 0.6044 | 0.4700 | 0.2845 | 0.1897 | 0.5429 | 0.3447 | 0.2926 | 0.1876 | 0.5328 | 0.3701 |
| IntSum | 0.3174 | 0.2304 | 0.6177 | 0.4826 | 0.2846 | 0.1965 | 0.5728 | 0.3521 | 0.2963 | 0.1979 | 0.5509 | 0.3714 |
| LLM-TRSR | 0.3143 | 0.2294 | 0.6189 | 0.4770 | 0.2822 | 0.1955 | 0.5781 | 0.3571 | 0.2954 | 0.1979 | 0.5525 | 0.3719 |
| CICL | 0.3267 | 0.2370 | 0.6233 | 0.4978 | 0.2868 | 0.1971 | 0.5684 | 0.3618 | 0.2980 | 0.1968 | 0.5510 | 0.3677 |
| Persona-DB | 0.3279 | 0.2430 | 0.6259 | 0.4839 | 0.2836 | 0.1956 | 0.5728 | 0.3589 | 0.2936 | 0.1939 | 0.5481 | 0.3669 |
| DPL (ours) | 0.3326 | 0.2445 | 0.6278 | 0.5136 | 0.2934 | 0.2011 | 0.5795 | 0.3812 | 0.3061 | 0.2061 | 0.5684 | 0.3780 |
4.1 Experimental Setup
Datasets. We focus on review generation, a representative personalized text generation task. Therefore, we conduct experiments on the Amazon Reviews 2023 dataset Hou et al. (2024)111https://amazon-reviews-2023.github.io/, focusing on the categories of Books, Movies & TV, and CDs & Vinyl. This dataset aggregates user-item interactions from Amazon, encompassing user reviews (ratings, texts, etc.) and item metadata (titles, descriptions, etc.). We preprocess the data into a format suitable for personalized review generation tasks, with details provided in Appendix A.
Baselines. We compare DPL with the following methods, with more details provided in Appendix B. Note that all methods are implemented using the same template and input format, as outlined in Appendix F, to ensure fair comparisons.
Non-Perso. This refers to the non-personalized approach, which completely excludes any user-specific information from the model input.
RAG Kumar et al. (2024). This is a method that employs a retrieval approach to extract key user-specific information, which is then used as instructional context for personalization.
IntSum Richardson et al. (2023). This method condenses the user’s retrieved history through summarization and incorporates it into the prompt. We achieve this by employing an LLM to generate a summary of the retrieved history.
LLM-TRSR Zheng et al. (2024). This method improves the IntSum method by segmenting the user’s history into blocks and employing a recurrent summarization paradigm to iteratively refine the summary within each new block.
CICL Gao and Das (2024). This Contrastive In-Context Learning method is not originally designed for LLM personalization. We adapt it to our experimental scenario by applying it to incorporate the review from another user into the prompt to an LLM summarizer, thereby improving the quality of summary generation.
Persona-DB Sun et al. (2025). This method introduces a collaborative refinement approach that allows a user to retrieve and integrate information from relevant peers. We apply it by incorporating the review of a relevant user into the prompt for the LLM generator.
Evaluation Metrics. We evaluate the experimental results using two complementary approaches, with additional details provided in Appendix C.
Conventional Evaluation. Following previous work on personalized text generation Kumar et al. (2024), we employ widely-adopted ROUGE-1 Lin (2004) and METEOR Banerjee and Lavie (2005) to measure lexical overlap between the generated reviews and ground-truth reviews, offering surface-level comparisons through n-gram matching and semantic alignment mechanisms Wang et al. (2023).
LLM-based Evaluation. Following the previous evaluation framework Salemi et al. (2025), we employ LLMs to assess the nuanced aspects of personalization, going beyond simple lexical matching. Specifically, the LLM evaluators process four key components – the generated review, the ground-truth review, review metadata, and item attributes – and produce raw scores ranging from 0 to 10. These scores are then normalized through min-max scaling to the [0, 1] range, with the evaluation metrics denoted as S-72B and S-GPT, corresponding to the use of Qwen-2.5-72B-Instruct-AWQ Yang et al. (2024) and GPT-4o-mini Achiam et al. (2023), respectively.
Implementation Details. For our DPL, we use BM25 Robertson and Zaragoza (2009) as the retriever and Qwen2.5-14B-Instruct Yang et al. (2024) as the LLM backbone, which functions as the difference extractor, summarizer, and review generator. For selective user comparison and structured difference extraction, we set the number of retrieved reviews, i.e., in Equation (3), to 8. Additionally, we utilize gte-Qwen2-1.5B-Instruct Li et al. (2023c) to map user reviews into embeddings, which are subsequently clustered into 5 groups by default, equivalent to setting the number of selected representative users in Equation (4) to 4. More details are provided in Appendix D.
4.2 Main Results (RQ1)
We begin by evaluating the overall performance of the compared methods. The summarized results are presented in Table 1, from which the following observations can be drawn:
DPL demonstrates superior personalization performance compared to the baseline methods across all metrics on three datasets. For instance, on the CDs & Vinyl dataset, it shows an average improvement of 2.84% over the best baseline metrics. This result underscores its enhanced personalized text generation capabilities, which can be attributed to its focus on extracting inter-user differences to improve LLM personalization through the selective comparison mechanism and difference extraction standard.
Non-Perso exhibits the weakest personalization performance across all experimental settings, which can be attributed to its fundamental limitation in incorporating user-specific contextual signals. In comparison, RAG demonstrates substantial performance gains by leveraging a retrieval approach to integrate user-specific information, thereby enabling more tailored outputs.
IntSum and LLM-TRSR demonstrate consistent superiority over RAG across all datasets, primarily due to their integration of summarization mechanisms. This is because summarization mechanisms distill the retrieved user history into more concise representations, thereby enhancing the utilization of contextual information.
Although CICL and Persona-DB demonstrate some performance improvements by utilizing the historical reviews of other relevant users, they still lag behind our DPL. This discrepancy arises because they fail to incorporate inter-user comparative analysis, which is essential for identifying the nuanced differences that genuinely influence user preferences.
4.3 Abalation Studies (RQ2)
To substantiate the rationale behind the various design decisions of DPL, we conduct an exhaustive evaluation using the Books dataset by systematically disabling one critical design element to obtain various variants. For the selective user comparison in Section 3.2.1, we replace the clustering-based selection method and introduce the following two variants for comparison:
Random. This variant randomly selects users without considering any specific criteria.
SimRank. This variant selects users based on the similarity of historical reviews between the current user and potential candidates. Specifically, the users with the top- lowest similarity are chosen in order to maximize differentiation.
For the structured difference extraction in Section 3.2.2, We modify the original extraction standard, whose dimensions include writing style, emotional style, and semantic style, and introduce the following four variants for comparison:
None. This variant excludes all the dimensions.
WriOnly, EmoOnly, SemOnly. These three variants each incorporate one of the three style dimensions, respectively.
| Methods | R-1 | MET. | S-72B | S-GPT |
| DPL | 0.3326 | 0.2445 | 0.6278 | 0.5136 |
| Random | 0.3282 | 0.2392 | 0.6246 | 0.5041 |
| SimRank | 0.3301 | 0.2421 | 0.6233 | 0.5078 |
| None | 0.3212 | 0.2362 | 0.6221 | 0.4877 |
| WriOnly | 0.3322 | 0.2461 | 0.6120 | 0.4959 |
| EmoOnly | 0.3241 | 0.2379 | 0.6151 | 0.5016 |
| SemOnly | 0.3244 | 0.2382 | 0.6196 | 0.5066 |
Table 2 illustrates the comparison results, from which we draw the following observations:
For the selective user comparison, replacing the clustering-based selection method results in a performance decline. This can be attributed to the fact that random user selection does not rely on any criteria, resulting in insufficient comparative information. Additionally, although SimRank introduces pairwise similarity for enhanced user selection, it fails to capture the global context offered by clustering, leading to a lack of representativeness in the selected users.
For the structured difference extraction, removing the parts of the complete extraction dimensions leads to a general performance decline. This finding suggests that each dimension plays a crucial role in capturing different facets of the differences, thereby ensuring meaningful and well-structured extraction outcomes for distinguishing task-relevant content.
4.4 In-depth Analyses (RQ3 & RQ4)
We next evaluate the performance of DPL across different values of hyperparameters and different levels of user uniqueness.
4.4.1 Hyper-parameter Analysis
In our investigation, the hyperparameter , which denotes the number of selected representative users for comparison, plays a pivotal role in determining the effectiveness of DPL. We vary within the set to assess the impact of varying these parameters on performance. Additionally, we present the best baseline metrics for comparative analysis, with the results shown in Figure 2. Our findings indicate that, generally, increasing the number of representative users leads to slight to moderate improvements in performance, with the best results occurring at . This underscores that effective difference extraction depends on selecting an appropriate number of comparative representatives.
4.4.2 Impact of User Uniqueness
We initially generate the uniqueness label based on users’ historical review data, as it does not inherently exist. Specifically, we map each user’s historical reviews to embeddings using LLMs and compute the average embedding of all users to establish a center. The Euclidean distance between each user’s embedding and this center is then calculated. Users with distances in the bottom 50% are classified as non-unique, while those in the top 50% are classified as unique.
Following this classification, we compare the performance of DPL with the non-personalization method across the two groups. For simplicity, we report the absolute improvement of R-1 and S-72B, as the trends for other metrics are consistent. As illustrated in Figure 3, the unique group shows a significantly greater improvement than the non-unique group. This suggests that the more unique the user, the better we can capture insights into their differences relative to others, aligning with the philosophy that “differences make us unique”.
4.5 Case Study (RQ5)
Finally, we conduct a case study to gain a deeper understanding of DPL’s capacity for personalized review generation. Specifically, we select a representative item from the Movies & TV dataset, and compare the reviews generated by DPL with those produced by the representative baseline method, LLM-TRSR. The detailed reviews are presented in Figure 4, where we highlight consistent sections between the generated reviews and the user’s real review by distinct colors.
We observe that the review generated by DPL exhibits more consistent sections. Furthermore, it offers a more personalized perspective and concise content, whereas the review generated by LLM-TRSR adopts a more objective stance and contains more tedious information. The alignment of DPL with the user’s writing style and semantic preferences further supports its superiority. More case studies are provided in Appendix G.
5 Related Work
LLM personalization has been explored across various domains, underscoring its importance. While existing works such as PUMA Cai et al. (2024) and PMG Shen et al. (2024) advance personalization in different scenarios, personalized text generation remains the foundation of LLM personalization Zhang et al. (2024b). LaMP Salemi et al. (2024b) evaluates short-form personalized text, later extended by LongLaMP Kumar et al. (2024) to assess long-form text generation. Among these, personalized item review writing stands out as a key task Peng et al. (2024) since user reviews naturally reflect individual preferences Ni et al. (2019), serving as a critical scenario to evaluate a model’s ability to generate personalized content.
Recent LLM-based enhance personalization using user-specific content, which can be categorized into: 1) retrieval-based methods, where ROPG Salemi et al. (2024a) introduces a retrieval selection model to dynamically choose the best retriever and HYDRA Zhuang et al. (2024) employs a personalized reranker to prioritize useful information; and 2) summarization-based methods, where ONCE Liu et al. (2024b) generates user profiles by summarizing topics and regions of interest in their browsing history and PPlug Liu et al. (2024a) converts user history into personalized embeddings.
Besides, a recent study Persona-DB Sun et al. (2025) has explored using data from other users to enhance personalized text generation. However, this method mainly focuses on enriching relevant knowledge rather than capturing each user’s uniqueness, which is crucial for effective personalization. To leverage inter-user data, contrastive learning Gao and Das (2024) offers a promising way of using both positive and negative examples to better describe intents. However, its performance is contingent on the identification of meaningful negative examples, which are challenging to obtain in user-centered LLM personalization. This reliance limits the full leverage of inter-user information for LLM personalization. To the best of our ability, our proposed DPL addresses this limitation.
6 Conclusions
This study introduced DPL, a novel method that enhances LLM personalization by identifying and leveraging inter-user differences as key preference information. We proposed a selective user comparison mechanism and a structured difference extraction method to extract task-relevant meaningful differences effectively. Experimental results on the personalized review generation task show that DPL significantly enhances LLM personalization. Currently, we only focus on review generation tasks to validate our method. In future, we plan to extend it to other tasks for broader verification. Additionally, our method’s effectiveness depends on the long-context processing capabilities of LLMs, and we currently assume that the context window is sufficient for handling our data. We also plan to extend our method to address long-context cases.
Limitations
In this paper, we utilize LLMs to extract inter-user differences, summarize difference-aware representations, and generate personalized reviews. These processes require a highly powerful model to effectively analyze users’ personalized information, ensuring that the generated reviews align closely with individual user preferences. Moreover, in this task, we specifically focus on three key aspects: writing style, emotional style, and semantic style. Our findings indicate that these factors significantly enhance the quality of personalized reviews. However, in other tasks, these standards may not be as suitable, and additional factors, such as personality traits, might need to be considered. Our approach involves comparing differences between users, which inevitably introduces additional computational overhead and time costs. We believe this challenge can be mitigated by precomputing and storing difference-aware representations in user memory. These representations can then be directly retrieved during online deployment, with periodic updates ensuring both efficiency and relevance.
Ethical Statements
Our method presents the following potential ethical concerns that need to be addressed carefully during deployment. First, it involves the storage and retrieval of users’ historical interaction data, which necessitates the implementation of strict safety protocols and ethical guidelines. Second, as we leverage inter-user differences to enhance personalization, sharing and integrating information between users raises significant privacy concerns. This process must incorporate robust privacy safeguards, including data anonymization, secure storage, and explicit user consent. Furthermore, the ability to create detailed user profiles for personalization purposes introduces the risk of potential misuse, such as surveillance or targeted manipulation. To mitigate these risks, developers bear the responsibility to prioritize privacy-preserving techniques and ensure that user data is handled with the utmost integrity and respect.
In our research, we utilize an open-source dataset, ensuring full compliance with the original dataset’s MIT license. Our research prioritizes data integrity, privacy, and ethical considerations to uphold the responsible and fair use of open-source resources.
Appendix A Dataset Details
In this paper, we focus on item review text generation, which is an effective task to evaluate the model’s ability to generate personalized content. To generate data samples, we employ the open-source Amazon Reviews Dataset Hou et al. (2024) and primarily follow the setting of LongLaMP Kumar et al. (2024), with several modifications to better align with our experimental framework.
Dataset Introduction. The Amazon Reviews dataset is processed by merging the user reviews subset with the item metadata subset using the asin field as the shared field. Specifically, the asin value from the user reviews subset is matched to the corresponding parent_asin field in the item metadata subset. From the user reviews subset, we extract asin, which is the unique identifier for the item; user_id, representing the unique identifier for the reviewer; title, which denotes the title of the review; text, containing the actual review content; rating, indicating the numerical rating given by the user; and timestamp, specifying the time when the review was submitted. From the item metadata subset, we select parent_asin, which serves as a reference to the asin field in the user reviews subset, linking reviews to items; title, representing the title of the item; and description, providing a textual summary of the item’s features. To distinguish between the review title and the item title, we use review_title for the former and item_title for the latter.
Data Curation. We perform a detailed data-cleaning process to create a high-quality benchmark. First, we ensure that each review includes a corresponding review_title, rating, and timestamp, with a minimum text length of 200 characters to meet the long-form criteria. For each item, we verify that it has an item_title and that its description falls within the 100 to 2000-character range to maintain high-quality content. Following this, we ensure that each product has at least four unique reviewers and retain only the most recent review from each reviewer for a given item. Additionally, we restrict the number of reviews per reviewer to between 18 and 500 to filter out users engaging in review manipulation. These settings are designed to align with our experiments.
Dataset Construction. After preprocessing, we proceed to construct the main experimental dataset. Here we follow the Temporal Setting used in LongLaMP, treating each user as a data point. However, unlike LongLaMP, we take a different approach when constructing the training dataset. Specifically, we use a user’s 10th most recent to 3rd most recent reviews as training samples, with their profile information derived from historical reviews relative to the target review. The validation dataset consists of the user’s 2nd most recent review, while the test dataset includes the 3rd most recent review. Each user’s historical information samples include only the item asin, rating, timestamp, review_title, and text. The asin references the pre-filtered item dataset, allowing us to extract the corresponding item information. Examples of the main experimental dataset and the item dataset are illustrated in Figure 5. We process the three categories: Books, Movies & TV, and CDs & Vinyl separately. The statistics for these categories after dataset construction are presented in Table 3.
Dataset Usage. In our experiments, we utilize only the test dataset for a direct evaluation of model performance. The details of model inputs are provided in Appendix F. Notably, the training and validation datasets are included to support related research.
| Categories () | #data | Profile Size | Output Length | |
| Books (839) | Train | 2536 | 29.3422.63 | 1295.77827.48 |
| Val | 317 | 33.8422.51 | 1216.49768.44 | |
| Test | 317 | 34.8422.51 | 1194.90801.18 | |
| Movies & TV (4832) | Train | 15400 | 35.6135.97 | 1691.421617.71 |
| Val | 1925 | 40.1135.89 | 1668.261612.17 | |
| Test | 1925 | 41.1135.89 | 1704.611751.98 | |
| CDs & Vinyl (3801) | Train | 14032 | 33.0032.44 | 1485.191292.09 |
| Val | 1754 | 37.5032.36 | 1602.931431.09 | |
| Test | 1754 | 38.5032.36 | 1600.041419.49 | |
Appendix B Baselines Details
| Method | Training-Free | Current User’s Data | Other Users’ Data | Difference-Aware |
| Non-Perso | ✓ | ✗ | ✗ | ✗ |
| RAG | ✓ | ✓ | ✗ | ✗ |
| IntSum | ✓ | ✓ | ✗ | ✗ |
| LLM-TRSR | ✓ | ✓ | ✗ | ✗ |
| CICL | ✓ | ✓ | ✓ | ✗ |
| Persona-DB | ✓ | ✓ | ✓ | ✗ |
| DPL | ✓ | ✓ | ✓ | ✓ |
In this section, we further introduce each baseline method in detail. Additionally, the comparison between different baseline methods and our proposed DPL is shown in Table 4.
Non-Perso:: This method does not utilize any user-specific information to assist in personalized review generation. The model’s input consists only of the item’s title and description, along with the rating and title of the review to be generated.
RAG: This method employs a retriever, specifically BM25 in our experiments, to retrieve relevant past reviews from the user’s review history. These retrieved reviews are then formatted and incorporated into the LLM’s input prompt.
IntSum: Unlike the original implementation of generating a summary offline and directly using it in downstream tasks, our scenario of long-form text review generation also involves long-form user history. Due to the input length limitations of LLMs, we cannot feed all of a user’s historical reviews into the model at once to produce a sensible user profile summary. Therefore, we explore an alternative solution: instead of generating a static summary beforehand, the model dynamically creates a summary based on the retrieved user’s historical reviews. This summary is then integrated into the model’s input prompt along with retrieved historical reviews. The experimental results demonstrate that our implementation is correct.
LLM-TRSR: To address the issue that our IntSum experiment does not utilize all of the user’s historical data, we select the LLM-TRSR method as a baseline. This method divides the user’s history into blocks and introduces two distinct preference summarization paradigms: hierarchical summarization and recurrent summarization. We opt for the recurrent summarization technique due to its superior performance and better suitability for the temporal setting of our dataset. In this method, a summary is first generated based on the initial history block and then iteratively updated using more recent history blocks. The final summary is then integrated into the model’s input along with the retrieved user’s historical reviews. In our experiments, each history block contains eight reviews.
CICL: This is the Contrastive In-Context Learning method. While originally not designed for LLM personalization, we adapt it to our experimental scenario. In our experiments, we treat the current user’s review as the positive example and a review from another user as the negative example. To construct these comparisons, we use the BM25 retriever to retrieve a historical review from the current user and randomly select a review on the same item from another user. These comparative reviews, along with the current user’s past reviews, are incorporated into the prompt for an LLM summarizer, thereby improving the quality of summary generation.
Persona-DB: This method introduces a collaborative refinement approach that allows a user to retrieve information from relevant peers to enrich knowledge. We adapt this approach to our task. For selecting relevant users, we employ an embedding-based strategy using the gte-Qwen2-1.5B-Instruct embedding model to encode users’ historical information and identify the most similar user to the current user. When generating the review, we first create a summary based on the current user’s retrieved historical reviews. Simultaneously, we extract the same number of reviews from the most similar user’s history. The current user’s profile summary, historical reviews, and the most similar user’s historical reviews are then integrated into the prompt for the LLM generator.
All baseline methods, including our approach, use prompts that are uniformly designed for our review generation task.
Appendix C Evaluation Metrics
| Datasets () | Books | Movies & TV | CDs & Vinyl | |||
| Methods () | R-L | BLEU | R-L | BLEU | R-L | BLEU |
| Non-Perso | 0.1438 | 1.2313 | 0.1262 | 0.3510 | 0.1262 | 0.4880 |
| RAG | 0.1617 | 4.1365 | 0.1414 | 1.7926 | 0.1413 | 2.0186 |
| IntSum | 0.1606 | 4.6217 | 0.1410 | 2.0803 | 0.1414 | 2.4589 |
| LLM-TRSR | 0.1616 | 5.0839 | 0.1408 | 2.1157 | 0.1416 | 2.4600 |
| CICL | 0.1654 | 5.3592 | 0.1422 | 2.1648 | 0.1412 | 2.3713 |
| Persona-DB | 0.1670 | 5.4166 | 0.1411 | 2.0175 | 0.1408 | 2.2963 |
| DPL | 0.1708 | 5.9593 | 0.1454 | 2.4065 | 0.1446 | 2.7514 |
| Methods () | R-1 | R-L | METEOR | BLEU | S-72B | S-GPT |
| Non-Perso | 0.3016 | 0.1508 | 0.1930 | 2.5779 | 0.6132 | 0.4650 |
| RAG | 0.3342 | 0.1765 | 0.2777 | 6.5834 | 0.6233 | 0.4833 |
| IntSum | 0.3252 | 0.1744 | 0.2851 | 6.4804 | 0.6278 | 0.4801 |
| LLM-TRSR | 0.3210 | 0.1718 | 0.2813 | 6.0386 | 0.6334 | 0.4890 |
| CICL | 0.3232 | 0.1679 | 0.2727 | 5.5111 | 0.6259 | 0.4890 |
| Persona-DB | 0.3252 | 0.1689 | 0.2768 | 5.6887 | 0.6183 | 0.4814 |
| DPL | 0.3366 | 0.1818 | 0.2879 | 6.6967 | 0.6461 | 0.5142 |
| Datasets () | Books | Movies & TV | CDs & Vinyl | ||||||||||
| Methods () | R-1 | MET. | S-72B | S-GPT | R-1 | MET. | S-72B | S-GPT | R-1 | MET. | S-72B | S-GPT | |
| Non-Perso | 0.2683 | 0.1512 | 0.5079 | 0.4360 | 0.2275 | 0.1304 | 0.4767 | 0.3242 | 0.2392 | 0.1336 | 0.4701 | 0.3409 | |
| RAG | N=1 | 0.3067 | 0.1820 | 0.5211 | 0.4114 | 0.2535 | 0.1533 | 0.4208 | 0.2476 | 0.2649 | 0.1553 | 0.4414 | 0.2989 |
| N=2 | 0.3167 | 0.1997 | 0.5609 | 0.4429 | 0.2682 | 0.1665 | 0.4712 | 0.2975 | 0.2755 | 0.1654 | 0.4803 | 0.3320 | |
| N=4 | 0.3220 | 0.2084 | 0.5987 | 0.4650 | 0.2785 | 0.1778 | 0.5117 | 0.3307 | 0.2880 | 0.1785 | 0.5158 | 0.3560 | |
| IntSum | N=1 | 0.3089 | 0.1950 | 0.5174 | 0.3628 | 0.2646 | 0.1677 | 0.3991 | 0.2079 | 0.2725 | 0.1653 | 0.3985 | 0.2391 |
| N=2 | 0.3172 | 0.2131 | 0.5703 | 0.4473 | 0.2758 | 0.1801 | 0.4798 | 0.2820 | 0.2854 | 0.1780 | 0.4888 | 0.3301 | |
| N=4 | 0.3188 | 0.2238 | 0.6076 | 0.4575 | 0.2821 | 0.1904 | 0.5368 | 0.3308 | 0.2932 | 0.1909 | 0.5364 | 0.3608 | |
| LLM- TRSR | N=1 | 0.3107 | 0.2024 | 0.5817 | 0.4442 | 0.2682 | 0.1747 | 0.5047 | 0.3018 | 0.2797 | 0.1747 | 0.5007 | 0.3431 |
| N=2 | 0.3146 | 0.2170 | 0.5975 | 0.4505 | 0.2749 | 0.1825 | 0.5289 | 0.3186 | 0.2883 | 0.1845 | 0.5271 | 0.3537 | |
| N=4 | 0.3174 | 0.2228 | 0.6164 | 0.4606 | 0.2803 | 0.1906 | 0.5557 | 0.3413 | 0.2931 | 0.1927 | 0.5413 | 0.3667 | |
| CICL | N=1 | 0.3157 | 0.2013 | 0.4921 | 0.3464 | 0.2652 | 0.1356 | 0.1689 | 0.2044 | 0.2741 | 0.1666 | 0.3921 | 0.2319 |
| N=2 | 0.3243 | 0.2162 | 0.5659 | 0.4221 | 0.2761 | 0.1798 | 0.4767 | 0.2804 | 0.2879 | 0.1805 | 0.4777 | 0.3092 | |
| N=4 | 0.3230 | 0.2282 | 0.5918 | 0.4669 | 0.2834 | 0.1889 | 0.5236 | 0.3271 | 0.2944 | 0.1805 | 0.5197 | 0.3488 | |
| Persona -DB | N=1 | 0.3184 | 0.2047 | 0.5281 | 0.4069 | 0.2667 | 0.1679 | 0.4243 | 0.2381 | 0.2731 | 0.1653 | 0.4115 | 0.2538 |
| N=2 | 0.3249 | 0.2228 | 0.5987 | 0.4631 | 0.2748 | 0.1773 | 0.4904 | 0.2984 | 0.2842 | 0.1773 | 0.4887 | 0.3239 | |
| N=4 | 0.3264 | 0.2305 | 0.6202 | 0.4902 | 0.2812 | 0.1882 | 0.5377 | 0.3321 | 0.2911 | 0.1867 | 0.5357 | 0.3623 | |
| DPL | N=1 | 0.3175 | 0.2010 | 0.5338 | 0.4069 | 0.2726 | 0.1731 | 0.4408 | 0.2542 | 0.2855 | 0.1761 | 0.4522 | 0.2829 |
| N=2 | 0.3256 | 0.2180 | 0.5823 | 0.4435 | 0.2858 | 0.1880 | 0.5117 | 0.3239 | 0.2995 | 0.1909 | 0.5162 | 0.3449 | |
| N=4 | 0.3270 | 0.2347 | 0.6095 | 0.4877 | 0.2896 | 0.1952 | 0.5493 | 0.3515 | 0.3059 | 0.2006 | 0.5532 | 0.3762 | |
In this paper, we consider the ROUGE-1, METEOR, S-72B, and S-GPT metrics in the primary experiments. We offer detailed descriptions of each metric as follows.
ROUGE-1: ROUGE-1 (Recall-Oriented Understudy for Gisting Evaluation) is a metric that measures the overlap of unigrams (individual words) between the generated text and the reference text.
METEOR: METEOR (Metric for Evaluation of Translation with Explicit ORdering) is a metric that evaluates text similarity by considering exact word matches, stemming, synonyms, and word order between the generated text and the reference text.
S-72B: S-72B is a metric that assesses the quality of generated text by comparing it to a reference text using the Qwen-2.5-72B-Instruct-AWQ model. We design the prompt following previous works Li et al. (2024c); Liu et al. (2023); Salemi et al. (2025) and make some reasonable modifications to better suit our task. In this evaluation, the prompt to the model includes the generated review, the user’s real review as the reference, as well as relevant metadata such as the review title, rating, and item details, including its title and description. We also perform scoring inference using vLLM, with the temperature set to 0.8 and top_p set to 0.95. The prompt is depicted in Figure C.
S-GPT: S-GPT is another metric that utilizes the GPT-4o-mini LLM as the evaluator. It adopts the same temperature, top_p, and prompt settings as S-72B. The final S-GPT score is calculated as the average of five evaluations.
Although traditional automated evaluation metrics such as ROUGE-L and BLEU Papineni et al. (2002) have been questioned for their ability to distinguish personalization from other subtle semantic aspects Koh et al. (2022), these metrics are still used in some studies Li et al. (2024a); Zhang et al. (2025a) and hold exploratory value. Therefore, we present a comparison of our method and the baseline methods based on ROUGE-L and BLEU, as shown in Table 5. The descriptions of these two metrics are provided below.
ROUGE-L: ROUGE-L is a metric that evaluates text similarity by measuring the longest common subsequence (LCS) between the generated text and the reference text.
BLEU: BLEU (Bilingual Evaluation Understudy) is a metric that evaluates text similarity by measuring the precision of n-gram matches between the generated text and the reference text. In our experiments, we use SacreBLEU Post (2018), which is a standardized version of BLEU. The higher the BLEU score, the more similar the generated text is to the reference.
Appendix D Implementation Details
D.1 Running Environments
We utilize the vLLM222https://github.com/vllm-project/vllm library Kwon et al. (2023) as the inference engine for LLMs in generating and the Sentence-Transformers333https://www.sbert.net/ Reimers and Gurevych (2019) library for LLMs in embedding in a Python 3.11.11 environment. All experiments are conducted on a single NVIDIA H100 NVL GPU with 96GB of GPU memory.
D.2 Hyperparameter Configurations
We configure the model with a maximum output length of 2048 tokens while keeping the input length unrestricted. Following the settings in LongLaMP, we set the temperature to 0.8 and the parameter top_p to 0.95. These settings remain consistent across all experiments.
Appendix E Additional Experiments
E.1 Model of Different Size
To investigate the generalization capability of our approach, we conduct experiments to evaluate its performance across models of different sizes. Specifically, we consider the Qwen2.5-32B-Instruct model. For each method, we utilize 8 retrieved historical data points. To reduce resource costs, we perform our experiments on the Books category only and report the results in ROUGE-1, ROUGE-L, METEOR, BLEU, S-72B, and S-GPT metrics. The experimental results are shown in Table 6, from which the following observations can be drawn:
When using the 32B model, its more powerful capabilities lead to performance improvements across almost all methods compared to the 14B model. However, DPL still achieves state-of-the-art performance, outperforming the best baseline metrics by an average of 2.26%. This demonstrates its strong generalization capability and effectiveness in the personalized review generation task.
CICL and Persona-DB show performance drops on certain metrics, with some even underperforming compared to RAG, IntSum, and LLM-TRSR. This suggests that simply leveraging other users’ information as auxiliary data is not particularly effective in enhancing the model’s personalization capabilities, emphasizing the idea that “differences make us unique”.
E.2 Number of Retrieved Reviews
We conducted experiments to explore the impact of different numbers of retrieved reviews on performance. Specifically, we evaluated cases with 1, 2, and 4 retrieved reviews. The experimental results presented in Table 7 lead to the following observations:
As the number of retrieved reviews increases, the performance metrics of all methods generally show an upward trend. This indicates that leveraging more historical user data enables the model to better infer users’ personalized preferences and generate more tailored reviews.
When the number of retrieved reviews is 4, our DPL method outperforms other baseline approaches on the CDs & Vinyl dataset and achieves competitive performance on the remaining two datasets. This demonstrates the superiority of our DPL method.
Appendix F Overview of Templates & Prompts
In this section, we provide the prompts used for different components. Specifically, we list the prompt used for the difference extractor LLM following Equation (5) in Figure 7, the prompt used for the summarizer LLM following Equation (3) in Figure 8, and the prompt used for the review generator LLM following Equation (2) in Figure 9.
Appendix G Case Studies
In this section, we provide more case studies to showcase our work.
G.1 Case Study: Comparison between Difference-aware Representation and User Profile Summary
We present a case study to compare the difference-aware representation generated by our proposed DPL method with the user profile summary produced by LLM-TRSR, as shown in Figure 10. From the figure, it is evident that the difference-aware representation generated by our method can better capture the user’s writing style, emotional style, and semantic style, thereby achieving a higher degree of personalization, as illustrated in Figure 4. In contrast, the profile summary produced by LLM-TRSR includes a lot of irrelevant information, which does not effectively contribute to personalized review generation.
G.2 Case Study: Outputs of DPL Components
We present a case study (Figure 11) to illustrate the outputs generated by the three key LLM-based components of DPL: the difference extractor, summarizer, and review generator. The results show that each component produces well-structured and meaningful outputs.